OpenStudy (anonymous):

The area of a right triangle is 260 in^2 Find the length of its legs if one leg is 6 inches longer than the other. Use the five step method

OpenStudy (mathstudent55):

I don't know what the five-step method is. This is how I'd solve this. In a right triangle, the legs form the right angle, so you can consider one leg to be the base and the other one to be the altitude in the formula for the area.

OpenStudy (mathstudent55):

\(A = \dfrac{bh}{2}\)

OpenStudy (anonymous):

Im almost done with it. I just need to solve L^2-6L-520=520

OpenStudy (anonymous):

**L^2-6L-520=0

OpenStudy (mathstudent55):

Let one leg be b. Then the other leg is 6 inches longer than b, or b + 6 Then \(260 = \dfrac{b(b + 6)}{2} \)

OpenStudy (mathstudent55):

\(b^2 + 6b - 520 = 0\)

OpenStudy (mathstudent55):

Ok, now you can try either factoring or the quadratic formula.

OpenStudy (anonymous):

b=20

OpenStudy (anonymous):

b=-26

OpenStudy (mathstudent55):

20 * 26 = 520 26 - 20 = 20

OpenStudy (mathstudent55):

\((b + 26)(b - 20) = 0\) b = 20 or b = -26 Since the side of a triangle can't have a negative length, we discard b = -26 One leg measures 20 in. The other leg is 6 inches longer, so it measures 26 in.

OpenStudy (anonymous):

u a life saver thanks fam!

OpenStudy (mathstudent55):

You're welcome.